OpenStudy (toxicsugar22):

is 18% of your grade alot or no

OpenStudy (watchmath):

no

OpenStudy (toxicsugar22):

what is it

OpenStudy (toxicsugar22):

mean

OpenStudy (watchmath):

If the perfect grade is 100. 18% of it is just 18 which is a very deep F.

OpenStudy (toxicsugar22):

so it would be 18 points is that it

OpenStudy (toxicsugar22):

so how much is 18% of your grade

OpenStudy (toxicsugar22):

how many points is that

OpenStudy (toxicsugar22):

if homework is 18% of your grade and u do homework evertime and get all of them right how much is that

OpenStudy (toxicsugar22):

canu someone help

OpenStudy (mattfeury):

think about it like a percentage. it is just math. 100% is a perfect score in the class. If you do great on homework and do it everyday and if it is correct (if it is graded for accuracy), then you will probably have 18 out of those 100 points towards your grade. if you don't do great on homework, say you have a 50% average on homework, then you will only have 9 points.

OpenStudy (anonymous):

what is your grade equals to ?

OpenStudy (toxicsugar22):

so out of thes tfour which ones are good homework 18% quizes 25% midterms 32% final 25%

OpenStudy (toxicsugar22):

so is 25 percent 25 points

OpenStudy (toxicsugar22):

im confused

OpenStudy (anonymous):

25% is 25 points of 100 that is equal to 25/100

OpenStudy (mattfeury):

yes all of those percentage can be considered points. also, all of your "points" on assignments / tests can be considered percentages.

OpenStudy (mattfeury):

if you get 10 out of 20 questions and they are all worth the same, then you got 10 / 20. 10 / 20 = 1 / 2 = .5 = 50% you would have a 50% on the assignment

OpenStudy (toxicsugar22):

so if we got 13/ 25 on a quiz what would that be

OpenStudy (mattfeury):

plug it in a calculator (you can just type it in google and it will calculate it) or you can do long division, but that's no fun :p

OpenStudy (toxicsugar22):

so if we do 13/ 2 and find out that percent and then if i got 25 32 and do that percent do i add thw two percentage and that is what i have as a grade so far

OpenStudy (toxicsugar22):

is that is

OpenStudy (toxicsugar22):

so if i do 20% on the first test and 60% on the second test do i add it up

OpenStudy (mattfeury):

take the average of the two tests and multiply by the percentage of your grade that tests are worth.

OpenStudy (toxicsugar22):

u mean as i said test are 32% of my grade

OpenStudy (toxicsugar22):

so i multiply it by that

OpenStudy (mattfeury):

yep

OpenStudy (anonymous):

18% is a reasonable amount

OpenStudy (toxicsugar22):

so if i had 20% on one test and 40% on the other the the total would be 60% divide b 2 which is 30 and multiply it by 32%

OpenStudy (anonymous):

not something you should just ignore

OpenStudy (toxicsugar22):

so if i had 20% on one test and 40% on the other the the total would be 60% divide b 2 which is 30 and multiply it by 32%

OpenStudy (toxicsugar22):

would that be it

OpenStudy (toxicsugar22):

is that how u do it

OpenStudy (anonymous):

mostly likely, I dont really understand you , what "mattfuery" typed seems correct

OpenStudy (mattfeury):

that is right. 30 * 32% = 30 * .32

OpenStudy (toxicsugar22):

so it would be 960

OpenStudy (toxicsugar22):

0r 9.6

OpenStudy (anonymous):

its pretty easy marks, to get 18% just for doing your homework, really doesnt get easier than that

OpenStudy (anonymous):

I know tons of people in university that do absolutely nothing all semester and just hope to get 50%'s on the finals (because they are worth 60% of the overall mark ) so they can pass

OpenStudy (anonymous):

you should always be looking to pick up easy marks

OpenStudy (toxicsugar22):

so 18 percet is alot

OpenStudy (toxicsugar22):

when it comes to your grade

OpenStudy (anonymous):

kind of depends how much your final exams are worth , if they worth 60% then its not "HEAPS" , if the finals are worth 20% then it would be considered alot

OpenStudy (anonymous):

I dont like to think of it that way though , it only takes a little application to pick up easy marks throughtout the year or semester

OpenStudy (anonymous):

Taking me as an example, I have only lost 1% out of the 35% already allocated to my maths course , the other 65% is on the final exam.

OpenStudy (anonymous):

35% of the course mark was based on fairly predictable , short exams (they had 3 or 4 questions and they were very similiar to the tutorial questions , no surprises at all )

OpenStudy (anonymous):

a little application can get alot of easy marks

OpenStudy (mattfeury):

think about it this way: what is worth the MOST when it comes to your grade? that should give you some perspective on how much is "a lot."

OpenStudy (toxicsugar22):

homework is 18% of the grade and final is 25% of the grade

OpenStudy (anonymous):

in that case it would be worth a fair bit

OpenStudy (anonymous):

ahh I miss the days of final exams worth 30% , all engineering uni exams are worth atleast 65% of final mark :| I only wish I could get more easy marks before the finals

OpenStudy (toxicsugar22):

so i am likely to pass if i did all my homework and i think i did good on the final

OpenStudy (mattfeury):

you are like to get close to 43% of your grade because that is 18% + 25%. it depends on the other sections. if you aced those two sections but didn't do anything else, then you would have a 43% which is an F, failing.

OpenStudy (anonymous):

its best not to worry bout marks , its a bit worry when people have to start calculating exactly what marks they need to pass

OpenStudy (toxicsugar22):

but i didd qizes and midterms

OpenStudy (anonymous):

be consistent , collect the easy marks along the way , put yourself in the best position to ace finals. The more work you do during the semester/year, the less you will have to do cramming before the final exam

OpenStudy (anonymous):

plus you get the easy marks

OpenStudy (toxicsugar22):

on midterms my scores were 30/100, 25/100, 27/100, 41/100

OpenStudy (toxicsugar22):

so is that good

OpenStudy (toxicsugar22):

will this bring my grade up

OpenStudy (mattfeury):

average them and multiply

OpenStudy (mattfeury):

but no those are not good marks :\. you didn't pass one of them.

OpenStudy (toxicsugar22):

i know but do i still get points

OpenStudy (anonymous):

yeh you still get "points" , but with those marks you will have to very lucky to get 50%

OpenStudy (mattfeury):

average the midterms. and multiply by the percentage midterms are worth. that will tell you how many points you get.

OpenStudy (toxicsugar22):

can u tell me how much they are i tried but i didnt get it

OpenStudy (anonymous):

:|

OpenStudy (mattfeury):

well what went wrong? this is why your grades are so low.

OpenStudy (anonymous):

not the best sign, when a maths student cant calculate their own weighted mark

OpenStudy (toxicsugar22):

so lets do this togethter so i know how to do it the nect time

OpenStudy (mattfeury):

do you know how to average numbers? add them up and divide by the number of terms. do you know how to add fractions? give them a common denominator. in this case they already have one (100). do you know how to multiple fractions? multiply your answer by the percentage midterms are worth. 32% = 32 / 100.

OpenStudy (mattfeury):

multiply*

OpenStudy (mattfeury):

let me know where you're having trouble.

OpenStudy (toxicsugar22):

ok i added the four test and i got 123/4 which is 30.75

OpenStudy (toxicsugar22):

did i do it right so far

OpenStudy (mattfeury):

the sum of yr tests is 123 / 100. so it is (123 / 100) / 4

OpenStudy (mattfeury):

when you add fractions, they must have the same denominator (bottom number). you add the numerators (top numbers) and put the result over the "common denominator."

OpenStudy (toxicsugar22):

can u show me once and leave it here

OpenStudy (toxicsugar22):

i am going to bed because it is midnight and i am tired

OpenStudy (toxicsugar22):

please

OpenStudy (mattfeury):

\[\frac{30}{100} + \frac{25}{100} + \frac{27}{100} + \frac{41}{100} = \frac{30 + 25 + 27 + 41}{100} = \frac{123}{100}\]

OpenStudy (mattfeury):

divide by the number of terms: \[\frac{\frac{123}{100}}{4} = \frac{123}{100} * \frac{1}{4} = \frac{123}{400}\]

OpenStudy (toxicsugar22):

yes i got thatt

OpenStudy (toxicsugar22):

so then what

OpenStudy (anonymous):

thats what u got, your current total , which is something like 27-28% , not good lols

OpenStudy (anonymous):

ohh no, its 30.7% , just checked on the calculator

OpenStudy (anonymous):

still , not very good considering you only have the final worth 25% left

OpenStudy (anonymous):

your absolutely best possible mark would be like 65% probably, I havent done the actual calculation but yeh

OpenStudy (toxicsugar22):

can u do the calculations please

OpenStudy (anonymous):

why, its probably best that you dont know what mark you need lols

OpenStudy (toxicsugar22):

so what does that mean